you need 25 ml 2% salt solution and you have 100 ml 5% salt solution. how much stock solution do you need? how much water?
Answer by lwsshak3(11628)

you need 25 ml 2% salt solution and you have 100 ml 5% salt solution. how much stock solution do you need? how much water?
**
let x=amt of 5% stock solution needed
25-x=water needed
5%x=2%*25
x=(2/5)*25=10
25-x=15
amt of 5% stock solution needed=10 ml
water needed=15 ml